OpenStudy (anonymous):

y= 3^x lnx can someone help me find the derivative please?

ganeshie8 (ganeshie8):

lookup product rule

OpenStudy (anonymous):

yes i understand that but I'm sort of confused on how to find the derivative of 3^x

jimthompson5910 (jim_thompson5910):

use this rule If \[\Large y = b^x\] then \[\Large \frac{dy}{dx} = b^x*\ln(b)\]

jimthompson5910 (jim_thompson5910):

notice that if b = e, then you'll have y = e^x become dy/dx = e^x*ln(e) = e^x

OpenStudy (anonymous):

so then it would be 3^x* ln(3) ?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

so would my final answer be 3^x/x + lnx * 3^x *ln(3) ?

jimthompson5910 (jim_thompson5910):

I think you mean \[\Large \frac{dy}{dx} = \frac{3^x}{x} + \ln(x)*3^x*\ln(3)\] right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

you are correct

OpenStudy (anonymous):

okay thank you :)

jimthompson5910 (jim_thompson5910):

np